# Asymptotic estimates for solutions of inhomogeneous non-divergence diffusion equations with drifts

### v1

@@ 1. Introduction
Equations (1.4), (1.5) and (1.6) fall into a larger class of equations which is formulated by (2.3). This nonlinear parabolic equation in non-divergence form will be the subject of our study. To the best of our knowledge, there are not many papers dealing with the long-time dynamics of non-divergence parabolic equations, especially the highly nonlinear ones such as (2.3). (However, see [14] for results for non-divergence linear elliptic problems with unbounded drift.) In our previous work [8], we essentially study (2.3) with zero forcing and, after shifting the solution by a constant, the homogeneous boundary condition. The purpose of the current paper is to obtain results for inhomogeneous problems. We emphasize that the equations we are dealing with contain drifts which can be nonlinear and unbounded as time tends to infinity. As we will see, it requires more technical and sophisticated tools.
@@ 2. Preliminaries
Throughout the paper, the spatial dimension $n\geq 1$ is fixed. For a vector $x\in\mathbb{R}^{n}$, its Euclidean norm is denoted by $|x|$. Let $\mathcal{M}^{n\times n}$ denote the set of $n\times n$ matrices of real numbers, and $\mathcal{M}^{n\times n}_{{\rm sym}}$ denote the set of symmetric matrices in $\mathcal{M}^{n\times n}$. For two matrices $A,B\in\mathcal{M}^{n\times n}$, their inner product $\langle A,B\rangle$ is the trace ${\rm Tr}(A^{\rm T}B)$. For a real-valued function $f(x)$ with $x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$, we denote by $D^{2}f$ the Hessian matrix of the second partial derivatives $(\partial^{2}f/\partial x_{i}\partial x_{j})_{i,j=1,\ldots,n}$.
@@
where $\mu$ is a continuous function from $J$ to $[0,\infty)$ or to $(-\infty,0]$. Indeed, in the former case, we just take $P$ such that $P^{\prime}=\mu$, while in the latter case, we take $P^{\prime}=-\mu$ and the new $K:=-K$. Then equation (2.7) is of the form (2.3) with the function $P$ satisfying (2.2).
@@ 3. Direct estimates by the Maximum Principle
()3.1 Define, for $t\in[0,T]$, $\mathcal{F}(t)=\int_{0}^{t}F(t){\rm d}\tau.$ Applying (i) and (ii) to $F_{1}=F$ and $F_{2}=F$, respectively, we have
@@
$u(\overline{U_{\infty}})\subset(-\infty,m_{1}+\widetilde{m}_{1}],\text{ resp.,}u(\overline{U_{\infty}})\subset[m_{2}-\widetilde{m}_{2},\infty).$ (3.7)
@@ 4. The linear problem
$A:U_{T}\to\mathcal{M}^{n\times n}_{{\rm sym}}\text{ and }b:U_{T}\to\mathbb{R}^{n}\text{ be given functions.}$. (4.1)
@@
Lemma 4.3 (Growth Lemma [8, Lemma 4.2]).
@@
$4c_{0}\beta_{k}\tau_{k}=R_{k}^{2}\text{ for all }k\geq 1$ (4.13)
Let $w\in C(\bar{U}\times[T_{0},\infty))\cap C_{x,t}^{2,1}(U\times(T_{0},\infty)$ and define
@@
Thanks to (4.15), we can apply, for each $k\geq 1$, Theorem 3.1, case ()3.1, part ()3.1 to the interval $[0,\tau_{k}]$ in place of $[0,T]$, functions $u(t):=w(T_{k-1}+t)$, $F_{1}(t):=F(T_{k-1}+t)$ and select $t=\tau_{k}$ in (3.2). It results in
@@
Let $k\geq 1$. We apply Lemma 4.4 to the interval $[T_{k-1},T_{k}]$ in place of $[0,T]$, with the same $M_{1}$, but, referring to the numbers in (4.11) and (4.12),
@@
Iterating this estimate for $k-1$, $k-2$, etc., we have
@@ 4.1. Case of bounded drifts
[…]
@@
[…]
@@
[…]
@@
Note that $\widetilde{L}w\geq-F(t)$, hence $\widetilde{L}(-w)\leq F(t).$ Applying part (i) to functions $(-w)$ and $F_{1}(t):=F(t)$, we obtain, recalling that $(-w)^{+}=w^{-}$,
@@
[…]
@@ 4.2. Case of unbounded drifts
    There is a continuous function $z_{0}(t)>0$ on $[T^{*},\infty)$ increasing to infinity as $t\to\infty$ such that
@@
    There is a decreasing, continuous function $\Lambda_{*}(t)\geq 0$ on $[T^{*},\infty)$ such that
@@
Referring to (4.35), we define a function $\bar{\Lambda}(t)$ on $[4T^{*},\infty)$ by
@@
$t\in[t_{*},\infty)\mapsto\bar{\Lambda}(t)\exp\left(e^{-\frac{d^{2}}{2c_{0}}-\varepsilon}\int_{t_{*}}^{t+1}e^{-\frac{d}{c_{0}}z_{0}(x)}{\rm d}x\right)\text{ is monotone.}$. (4.38)
@@
$-\ln(1-x)=x+\frac{1}{2}x^{2}+\mathcal{O}(x^{3})\text{ as }x\to 0.$
@@
Step 3. We will apply Proposition 4.5. Let $k\geq 1$. Noticing from (4.50) that $[T_{k-1},T_{k}]\subset[T_{0},T_{0}+k]$, hence, we have on $U\times(T_{k-1},T_{k}]$
@@
Because $\Lambda_{*}(t)$ is decreasing and $03T_{0}-1>$ 0, we have for $k\geq 1$,
@@
[…]
@@
Because $z(j)$ is increasing in $j$, so is $-e^{-z(j)}$ and, hence, we can estimate
$-e^{-z(j)}\leq-\int_{j}^{j+1}e^{-z(x)}{\rm d}x.$
@@
[…]
@@
where $\varphi(x)=x$ and $h(t)=e^{-z(t)}$. Note from condition (4.37) that $\int_{0}^{\infty}h(\tau){\rm d}\tau=\infty.$ By applying [9, Lemma A.1] to equation (4.69), we obtain
@@
We apply Theorem 4.11 to $w$ and $(-w)$, and then use inequality (4.29) to obtain (4.73). We omit the details. ∎
@@
If $\alpha\in(0,1)$, then the term $-(1+\gamma\delta)/t$ in (4.77) dominates as $t\to\infty$, hence $\mathcal{F}^{\prime}(t)<0$ for large $t$.
@@ 5.1. Reduction method
The following lemma is an extension of [8, Lemma 3.3] to inhomogeneous problems.
@@
We obtain from [8, Eq. (3.18)] that
@@
We recall the following examples from [8, Example 3.4].
@@
    $F_{\lambda}(s)=\frac{1}{\lambda}e^{\lambda s}\text{ for all }s\in J,\text{ or }F_{\lambda}(s)={\rm sign}(\lambda)e^{\lambda s}\text{ for all }s\in J.$ (5.11)
@@
    $F_{\lambda}(s)=\frac{1}{\lambda+1}s^{\lambda+1}\text{ for all }s>0,\text{ or }F_{\lambda}(s)={\rm sign}(\lambda+1)s^{\lambda+1}\text{ for all }s>0.$ (5.12)
@@ 5.2. Convergence of the solution
    $-c_{1}|\xi|^{2}\leq\xi^{\rm T}K(x,t)\xi\leq c_{2}|\xi|^{2}\text{ for all $(x,t)\in U_{\infty}$, all $\xi\in\mathbb{R}^{n}$.}$ (5.13)
@@
$u(\overline{U_{\infty}})\subset[m_{*},M_{*}]\text{ for some numbers $m_{*},M_{*}\in J$.}$ (5.20)
@@
Using the first inequality in (5.26), we can apply Theorem 4.7()4.7 to operator $\widetilde{L}:=\mathcal{L}$, function $w:=\bar{w}_{1}$, number $M_{2}:=\mu_{4}$ and function $F_{1}(t):=\mathcal{F}(t)$. We obtain from (4.21) that
@@
Similarly, using the second inequality in (5.26), we can apply Theorem 4.7()4.7 to operator $\widetilde{L}=\mathcal{L}$, function $w:=-\bar{w}_{2}$, number $M_{2}=\mu_{4}$ and function $F_{1}(t):=\mathcal{F}(t)$. We obtain from (4.21) that
@@
[…]
@@
    There are numbers $u_{*}\in J$, $T^{*}>0$ and a decreasing, continuous function $\widetilde{\Lambda}(t)\geq 0$ on $[T^{*},\infty)$ such that
@@
that is, we obtain inequality (5.31) again.
@@
Thus, we obtain inequality (5.32) again.
@@
$u(\overline{U_{\infty}})\subset[m,\infty)\text{ for some number $m>0$,}$ (5.46)
@@
$|u(x,t)|\leq\mu_{1}\mathcal{F}(t),\text{ where $\mu_{1}$ is some positive number.}$. (5.52)
@@
Step 2. Let $\lambda_{2}<0$ such that $-1\neq\lambda_{2}\leq-c_{2}/c_{0}$. For $j=1,2$, define the function $w_{j}=F_{\lambda_{j}}(u):=\frac{1}{\lambda_{j}+1}u^{\lambda_{j}+1}$ on $\overline{U_{\infty}}$, see the first formula in (5.12) of Example 5.2()5.2. Let $\bar{u}=u-u_{*}$, and, for $j=1,2$, let $w_{*,j}=F_{\lambda_{j}}(u_{*})$ and define $\bar{w}_{j}=w_{j}-w_{*,j}$ on $\overline{U_{\infty}}$. Since property (5.50) implies (5.17), we obtain inequality (5.27) again, i.e., for all $x\in\Gamma$ and sufficiently large $t$,
@@
By (5.54) and the definition of $\mathcal{F}_{1}(t)$ in (5.56), we have, for large $t$,
@@
From this we find that, see e.g. [8, Inequalities (4.60) and (4.61)],
@@
Consider sufficiently large $t$. By (5.48), we have $m^{\lambda_{2}}\leq\mathcal{F}(t)^{\lambda_{1}}$. Together with (5.54), this yields
@@
$-\varepsilon\leq\min_{x\in\bar{U}}\bar{u}(x,t)\leq\bar{u}(x,t)\leq\max_{x\in\bar{U}}\bar{u}(x,t)\leq\varepsilon\text{ for all }x\in\bar{U},$
@@
We follow the proof of Theorem 5.9. In this case $g\geq 0$, $u\geq 0$ and $u_{*}=0$. Although $P(s)$ is not defined at $s=u_{*}=0$, we have for $\lambda_{1}>0$ the transformation $F_{\lambda_{1}}(s)=\frac{s^{\lambda_{1}+1}}{\lambda_{1}+1}$ for $s>0$ can be extended to the function $\widehat{F}(s)=\frac{s^{\lambda_{1}+1}}{\lambda_{1}+1}$ for all $s\geq 0$. Note that $\widehat{F}^{\prime}(s)=s^{\lambda_{1}}$ is bounded near 00. In the proof of Theorem 5.9, we replace $F_{\lambda_{1}}$ with $\widehat{F}$ and proceed up to (5.58). In this case $\bar{w}_{1}=w_{1}=\widehat{F}(u)\geq 0$. Then inequality (5.58) reads as
@@
    For more cases of $u(\overline{U_{\infty}})\subset\bar{J}$ and/or $u_{*}\in\partial J\setminus J$, see [8, Theorem 4.7].
@@
Funding. No funds were received for conducting this study.

### v2

@@ 1. Introduction
Equations (1.4), (1.5) and (1.6) fall into a larger class of equations which is formulated by (2.3) below. This nonlinear parabolic equation in non-divergence form will be the subject of our study. To the best of our knowledge, there are not many papers dealing with the long-time dynamics of non-divergence parabolic equations, especially the highly nonlinear ones such as (2.3). See [12, 13] for the general existence and regularity theory. See also [16] for results for non-divergence linear elliptic problems with unbounded drift. In our previous work [9], we essentially study (2.3) with zero forcing and, after shifting the solution by a constant, the homogeneous boundary condition. The purpose of the current paper is to obtain results for inhomogeneous problems. We emphasize that the equations we are dealing with contain drifts which can be nonlinear and unbounded as time tends to infinity. As we will see, it requires more technical and sophisticated tools.
@@ 2. Preliminaries
Throughout the paper, the spatial dimension $n\geq 1$ is fixed. For a vector $x\in\mathbb{R}^{n}$, its Euclidean norm is denoted by $|x|$. Let $\mathcal{M}^{n\times n}$ denote the set of $n\times n$ matrices of real numbers, and $\mathcal{M}^{n\times n}_{{\rm sym}}$ denote the set of symmetric matrices in $\mathcal{M}^{n\times n}$. For two matrices $A,B\in\mathcal{M}^{n\times n}$, their inner product $\langle A,B\rangle$ is the trace ${\rm Tr}(A^{\rm T}B)$. For a scalar function $f(x)$ with $x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$, we denote by $D^{2}f$ the Hessian matrix of the second partial derivatives $(\partial^{2}f/\partial x_{i}\partial x_{j})_{i,j=1,\ldots,n}$.
@@
where $\mu$ is a continuous function from $J$ to $0,\infty)$ or to $(-\infty,0]$. Indeed, in the former case, we just take $P$ such that $P^{\prime}=\mu$, while in the latter case, we take $P^{\prime}=-\mu$ and the new K:=KK:=-K. Then equation (2.7) is of the form (2.3) with the function $P$ satisfying (2.2).
@@ 3. Direct estimates by the Maximum Principle
()3.1 Define, for $t\in[0,T]$, $\mathcal{F}(t)=\int_{0}^{t}F(t){\rm d}\tau.$ Applying (i) and ($i$) to $F_{1}=F$ and $F_{2}=F$, respectively, we have
@@
[…]
@@ 4. The linear problem
$A:U_{T}\to\mathcal{M}^{n\times n}_{{\rm sym}}\text{ and }b:U_{T}\to\mathbb{R}^{n}\text{ be given functions.}$ (4.1)
@@
Lemma 4.3 (Growth Lemma [9, Lemma 4.2]).
@@
$4c_{0}\beta_{k}\tau_{k}=R_{k}^{2}\text{ for all }k\geq 1.$ (4.13)
Let $w\in C(\bar{U}\times[T_{0},\infty))\cap C_{x,t}^{2,1}(U\times(T_{0},\infty))$ and define
@@
Thanks to (4.15), we can apply, for each $k\geq 1$, Theorem 3.1, case ()3.1, part ()3.1 to the interval $0,\tau_{k}]$ in place of $[0,T]$, functions $u(t):=w(T_{k-1}+t)$, $F_{1}(t):=F(T_{k-1}+t)$ and select $t=\tau_{k}$ in (3.2). It results in
@@
Let $k\geq 1$. We apply Lemma 4.4 to the interval $T_{k-1},T_{k}]$ in place of $[0,T]$, with the same $M_{1}$, but, referring to the numbers in (4.11) and (4.12),
@@
Iterating this estimate for k1k-1, k2k-2, etc., we have
@@ 4.1. Case of bounded drifts
$\limsup_{k\to\infty}J_{k}\leq\limsup_{k\to\infty}\left(\sum_{m=0}^{k}\eta_{*}^{k-m}\Lambda_{m}\right)\leq\frac{1}{1-\eta_{*}}\limsup_{k\to\infty}\Lambda_{k}.$ (4.24)
@@
$\limsup_{t\to\infty}\left(\max_{x\in\bar{U}}w^{+}(x,t)\right)\leq\limsup_{k\to\infty}J_{k-1}+\limsup_{k\to\infty}\Lambda_{k}\leq\frac{2-\eta_{*}}{1-\eta_{*}}\limsup_{k\to\infty}\Lambda_{k}.$ (4.25)
@@
$\limsup_{k\to\infty}\Lambda_{k}\leq\limsup_{k\to\infty}\delta_{k}+\limsup_{k\to\infty}D_{k}\leq\limsup_{t\to\infty}\left(\max_{x\in\Gamma}w^{+}(x,t)\right)+\limsup_{t\to\infty}\int_{t}^{t+T_{*}}F_{1}(\tau){\rm d}\tau.$ (4.26)
@@
Note that $\widetilde{L}w\geq-F(t)$, hence $\widetilde{L}(-w)\leq F(t).$ Applying part (i) to functions (w)(-w) and $F_{1}(t):=F(t)$, we obtain, recalling that $(-w)^{+}=w^{-}$,
@@
$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}|w(x,t)|\right]\leq\max\left\{\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}w^{+}(x,t)\right],\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}w^{-}(x,t)\right]\right\}.$ (4.29)
@@ 4.2. Case of unbounded drifts
    There is a continuous function $z_{0}(t)>0$ on $T^{*},\infty)$ increasing to infinity as $t\to\infty$ such that
@@
    There is a decreasing, continuous function $\Lambda_{*}(t)\geq 0$ on $T^{*},\infty)$ such that
@@
Referring to (4.35), we define a function $\bar{\Lambda}(t)$ on $4T^{*},\infty)$ by
@@
$t\in[t_{*},\infty)\mapsto\bar{\Lambda}(t)\exp\left(e^{-\frac{d^{2}}{2c_{0}}-\varepsilon}\int_{t_{*}}^{t+1}e^{-\frac{d}{c_{0}}z_{0}(x)}{\rm d}x\right)\text{ is monotone.}$ (4.38)
@@
$\ln(1-x)=x+\frac{1}{2}x^{2}+\mathcal{O}(x^{3})\text{ as }x\to 0.$
@@
Step 3. We will apply Proposition 4.5. Let $k\geq 1$. Noticing from (4.50) that $T_{k-1},T_{k}]\subset[T_{0},T_{0}+k]$, hence, we have on $U\times(T_{k-1},T_{k}]$
@@
For $k\geq 1$, estimating $T_{k-1}$ by the first inequality in (4.50), and using the facts that $\Lambda_{*}(t)$ is decreasing and $3T_{0}-1>0$, we have
@@
$\limsup_{t\to\infty}\left(\max_{x\in\bar{U}}w^{+}(x,t)\right)\leq\limsup_{k\to\infty}J_{k-1}+\limsup_{k\to\infty}\Lambda_{k}\leq\limsup_{k\to\infty}J_{k}.$ (4.58)
@@
Because $z(j)$ is increasing in $j$, so is $e^{-z(j)}$ and, hence, we can estimate
$e^{-z(j)}\leq-\int_{j}^{j+1}e^{-z(x)}{\rm d}x.$
@@
$\limsup_{k\to\infty}J_{k}\leq\limsup_{k\to\infty}\sum_{m=0}^{k}I_{k,m}\Lambda_{m}\leq\limsup_{k\to\infty}\widetilde{J}_{k},$ (4.61)
@@
where $\varphi(x)=x$ and $h(t)=e^{-z(t)}$. Note from condition (4.37) that $\int_{0}^{\infty}h(\tau){\rm d}\tau=\infty.$ By applying [8, Lemma A.1] to equation (4.69), we obtain
@@
We apply Theorem 4.11 to $w$ and (w)(-w), and then use inequality (4.29) to obtain (4.73). We omit the details. ∎
@@
If $\alpha\in(0,1)$, then the term $(1+\gamma\delta)/t$ in (4.77) dominates as $t\to\infty$, hence $\mathcal{F}^{\prime}(t)<0$ for large $t$.
@@ 5.1. Reduction method
The following lemma is an extension of [9, Lemma 3.3] to inhomogeneous problems.
@@
We obtain from [9, Eq. (3.18)] that
@@
We recall the following examples from [9, Example 3.4].
@@
[…]
@@
[…]
@@ 5.2. Convergence of the solution
    $c_{1}|\xi|^{2}\leq\xi^{\rm T}K(x,t)\xi\leq c_{2}|\xi|^{2}\text{ for all $(x,t)\in U_{\infty}$, all $\xi\in\mathbb{R}^{n}$.}$ (5.13)
@@
[…]
@@
Using the first inequality in (5.26), we can apply Theorem 4.7()4.7 to operator $\widetilde{L}$$\mathcal{L}$, function $w:=\bar{w}_{1}$, number $M_{2}:=\mu_{4}$ and function $F_{1}(t):=\mathcal{F}(t)$. We obtain from (4.21) that
@@
Similarly, using the second inequality in (5.26), we can apply Theorem 4.7()4.7 to operator $\widetilde{L}$$\mathcal{L}$, function $w:=-\bar{w}_{2}$, number $M_{2}=\mu_{4}$ and function $F_{1}(t):=\mathcal{F}(t)$. We obtain from (4.21) that
@@
$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}|\bar{u}(x,t)|\right]\leq\max\left\{\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{u}^{+}(x,t)\right],\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{u}^{-}(x,t)\right]\right\}$
@@
    There are numbers $u_{*}\in J$, $T^{*}>0$ and a decreasing, continuous function $\widetilde{\Lambda}(t)\geq 0$ on $T^{*},\infty)$ such that
@@
that is, we obtain (5.31) again.
@@
Thus, we obtain (5.32) again.
@@
[…]
@@
$|u(x,t)|\leq\mu_{1}\mathcal{F}(t),\text{ where $\mu_{1}$ is some positive number.}$ (5.52)
@@
Step 2. Let $\lambda_{2}<0$ such that $1\neq\lambda_{2}\leq-c_{2}/c_{0}$. For $j=1,2$, define the function $w_{j}=F_{\lambda_{j}}(u):=\frac{1}{\lambda_{j}+1}u^{\lambda_{j}+1}$ on $\overline{U_{\infty}}$, see the first formula in (5.12) of Example 5.2()5.2. Let $\bar{u}=u-u_{*}$, and, for $j=1,2$, let $w_{*,j}=F_{\lambda_{j}}(u_{*})$ and define $\bar{w}_{j}=w_{j}-w_{*,j}$ on $\overline{U_{\infty}}$. Since property (5.50) implies (5.17), we obtain inequality (5.27) again, i.e., for all $x\in\Gamma$ and sufficiently large $t$,
@@
By (5.54), the definition of $\mathcal{F}_{1}(t)$ in (5.56), and inequality (5.45) we have, for large $t$,
@@
From this we find that, see e.g. [9, Inequalities (4.60) and (4.61)],
@@
Consider sufficiently large $t$. By (5.48), we have $m^{\lambda_{2}}\leq\mathcal{F}(t)^{\lambda_{1}}$. Together with (5.54) and (5.45), this yields
@@
$\varepsilon\leq\min_{x\in\bar{U}}\bar{u}(x,t)\leq\bar{u}(x,t)\leq\max_{x\in\bar{U}}\bar{u}(x,t)\leq\varepsilon\text{ for all }x\in\bar{U},$
@@
We follow the proof of Theorem 5.9. In this case $g\geq 0$, $u\geq 0$ and $u_{*}=0$. Although $P(s)$ is not defined at $s=u_{*}=0$, we have for $\lambda_{1}>0$ the transformation $F_{\lambda_{1}}(s)=\frac{s^{\lambda_{1}+1}}{\lambda_{1}+1}$ for $s>0$ can be extended to the function $\widehat{F}(s)=\frac{s^{\lambda_{1}+1}}{\lambda_{1}+1}$ for all $s\geq 0$. Note that $\widehat{F}^{\prime}(s)=s^{\lambda_{1}}$ is bounded near 0. In the proof of Theorem 5.9, we replace $F_{\lambda_{1}}$ with $\widehat{F}$ and proceed up to (5.58). In this case $\bar{w}_{1}=w_{1}=\widehat{F}(u)\geq 0$. Then inequality (5.58) reads as
@@
    For more cases of $u(\overline{U_{\infty}})\subset\bar{J}$ and/or $u_{*}\in\partial J\setminus J$, see [9, Theorem 4.7].
@@
Funding. A.I. obligatorily acknowledges OGRI’s grant 122022800272-4.